# WritingBench: A Comprehensive Benchmark for Generative Writing

### v3

@@ Abstract
Recent advancements in large language models (LLMs) have significantly enhanced text generation capabilities, yet evaluating their performance in generative writing remains a challenge. Existing benchmarks primarily focus on generic text generation or limited in writing tasks, failing to capture the diverse requirements of high-quality written contents across various domains. To bridge this gap, we present WritingBench, a comprehensive benchmark designed to evaluate LLMs across 6 core writing domains and 100 subdomains, encompassing creative, persuasive, informative, and technical writing. We further propose a query-dependent evaluation framework that empowers LLMs to dynamically generate instance-specific assessment criteria. This framework is complemented by a fine-tuned critic model for criteria-aware scoring, enabling evaluations in style, format and length. The framework’s validity is further demonstrated by its data curation capability, which enables 7B-parameter models to approach state-of-the-art (SOTA) performance. We open-source the benchmark, along with evaluation tools and modular framework components, to advance the development of LLMs in writing.
WritingBench: A Comprehensive Benchmark for Generative Writing
Yuning Wu1, Jiahao Mei1,3, Ming Yan1thanks: Corresponding authors, Chenliang Li1, Shaopeng Lai1, Yuran Ren2, Zijia Wang2, Ji Zhang1, Mengyue Wu3, Qin Jin2*, Fei Huang1 1Alibaba Group, 2Renmin University of China, 3Shanghai Jiao Tong University {yuningwu, ym119608}@alibaba-inc.com, qjin@ruc.edu.cn
@@ 1 Introduction
Figure 1: WritingBench query example with color-coded requirements. The three black-bordered categories highlight essential requirements analyzed in follow-up assessments. Red phrases correlate with gray-shaded writing support materials.
@@
In recent years, LLMs DeepSeek-AI et al. (2025); Hurst et al. (2024) have revolutionized text generation, demonstrating impressive performance across diverse applications, from generating creative content Mirowski et al. (2023); Lei et al. (2024) and assisting in education Wang et al. (2024a); Li et al. (2024b) to enhancing professional workflows Shao et al. (2024); Li et al. (2024a). However, generative writing, which requires high levels of coherence, creativity, logical reasoning, and stylistic precision, poses a unique challenge that existing evaluation methods fail to address adequately.
@@
Current evaluation benchmarks for generative writing suffer from two major limitations: 1) Limited scope and diversity in task formulation; and 2) Inadequate evaluation metrics for complex writing tasks. First, there is a significant lack of specialized benchmarks covering a broad range of writing tasks. Most existing writing-oriented benchmarks are restricted to single domains, such as fictions Karpinska et al. (2024); Gómez-Rodríguez and Williams (2023), and their task formulations tend to be simplistic—often relying on single-sentence queries Bai et al. (2024) or a small set of instruction templates Paech (2023); Que et al. (2024). Additionally, many benchmarks use homogeneous input materials Que et al. (2024); Karpinska et al. (2024), limiting their ability to accommodate the complex and customized requirements inherent in real-world writing scenarios. As a result, they fail to capture the diversity and intricacies of practical writing tasks (see Figure 1). Second, current automatic evaluation metrics lack the robustness needed for a comprehensive and nuanced assessment of writing quality. While LLM-based evaluation methods show promise in capturing semantic meanings Shao et al. (2024); Que et al. (2024); Bai et al. (2024), they typically rely on a narrow set of predefined criteria (e.g., fluency and coherence). As LLMs continue to evolve with increasingly sophisticated writing capabilities, these static evaluation criteria and frameworks are inadequate for assessing the complex, multi-dimensional nature of writing, including creativity, argumentation strength, and domain-specific adherence.
@@
To address these challenges, we introduce WritingBench, a comprehensive benchmark and robust framework for evaluating general-purpose writing. Our approach begins with a carefully designed secondary domain categorization, grounded in real-world writing needs. We develop a four-stage query construction pipeline (illustrated in Figure 2), where LLMs first generate and diversify writing queries, followed by human-driven material collection and optimization. This process ensures a diverse set of writing tasks with broad domain coverage, varied requirements, and integration of heterogeneous source of materials. To enable a more nuanced evaluation, we propose a query-dependent evaluation framework that dynamically generates five instance-specific criteria using LLMs, which are then scored by a fine-tuned critic model. Finally, we integrate the framework to filter writing-specific data and train a small-scale model to verify its ability in identifying high-quality writing samples.
Our primary contributions are as follows:
@@
$\bullet$ We present WritingBench, an open-source writing benchmark comprising 1,239 queries across 6 primary domains and 100 subdomains,featuring style, format and length requirements. WritingBench supports extended-context generation with input ranging from tens to thousands of words, addressing real-world diversity. It facilitates systematic evaluation to identify improvement areas and highlights the potential of chain-of-thought (CoT) processes in creative tasks.
@@
$\bullet$ We propose a query-dependent evaluation framework that integrates instance-specific criteria generation with a criteria-aware scoring model. It achieves 83% human alignment, significantly surpassing static-criteria baselines (65%, 59%). The effectiveness is further evidenced by its data curation capability - models trained with framework-filtered data match SOTA performance.
@@
$\bullet$ We publicly release WritingBench, including its evaluation protocols, criteria generation tools with an integrated critic model, and writing-enhanced models, to foster further research. Available at: https://github.com/X-PLUG/WritingBench.
@@ 2.1 Writing Benchmarks
Existing evaluation benchmarks suffer from significant limitations in domain coverage and task granularity. For instance, EQ-Bench encompasses story-related quires using templated queries Paech (2023), while LongBench-Write incorporates length constraints in 120 queries Bai et al. (2024); however, they both lack hierarchical domain taxonomies and multi-dimensional requirement specifications (e.g., style and format). Furthermore, most current benchmarks rely on fixed instruction templates, short contexts or materials predominantly from a single source Que et al. (2024); Karpinska et al. (2024); Marco et al. (2024), rendering them insufficient for addressing the complexity of real-world data needs. In contrast, our proposed benchmark fills these gaps by introducing 1,239 free-form queries distributed across 6 primary domains and 100 subdomains, with potential controls over style, format, and length, paired with inputs ranging from tens to thousands of words.
@@
WritingBench (Ours) 1,239 6 100 1,546 19,361
Table 1: Comparison of existing writing benchmarks.
@@ 2.2 Evaluation Methods
Using LLMs as evaluators has become a prevalent approach for evaluating the quality of generated responses. Typically, researchers pre-define a fixed set of evaluation dimensions applicable across all test instances Gómez-Rodríguez and Williams (2023); Marco et al. (2024). For example, SuperCLUE Xu et al. (2023) employs three dimensions, whereas LongBench-Write Bai et al. (2024) adopts six dimensions (e.g., relevance and accuracy). HelloBench Que et al. (2024) introduces task-specific dimensions, but the dimensions remain consistent across all queries of a given task. Although the LLM-as-a-judge approach enhances scalability, static evaluation dimensions often fail to accommodate the diversity of writing styles and specifications. To address this limitation, recent work Liang et al. (2024) trains a model to dynamically generate evaluation dimensions for individual queries. However, the dimensions remains confined to a small predefined set. In contrast, our query-dependent evaluation framework leverages LLMs to generate diverse and instance-specific criteria while fine-tuning a dedicated critic model to perform the evaluation.
@@ 2.3 Writing-Enhanced Models
Although existing LLMs demonstrate exceptional writing capabilities, researchers strive to further enhance their overall writing proficiency. Recent models, such as Weaver Wang et al. (2024b) and Suri Pham et al. (2024), have exhibited notable domain-specific strengths. For instance, Weaver benefits from over 200B parameter pretraining, supporting four distinct writing domains, while LongWriter Bai et al. (2024) specializes in length constraints. However, these models experience substantial performance degradation in cross-domain scenarios and multi-constraint tasks. In this work, leveraging the effectiveness of our evaluation framework, we introduce writing-enhanced models trained on high-quality data, achieving high performance across various tasks.
@@ 3 WritingBench
Figure 3: Domain categories in WritingBench. Inner represents the 6 primary domains and outer depicts 100 secondary subdomains (Sub = subdomains per category, Num = dataset entries).
@@
Academic & Engineering 187 1,915 15,534
Finance & Business 238 1,762 19,361
Politics & Law 226 2,274 18,317
Literature & Arts 242 1,133 9,973
Education 151 1,173 10,737
Advertising & Marketing 195 886 6,504
@@
Style 395 1,404 18,197
Format 342 1,591 18,197
Length 214 1,226 14,097
@@
<1K 727 443 994
1K-3K 341 1,808 2,991
3K-5K 94 3,804 4,966
5K+ 77 8,042 19,361
Table 2: Data statistics for WritingBench categorized by domain, requirement, and length.
In this section, we will mainly introduce the construction process of our WritingBench, and the query-dependent evaluation framework with the accompanied critic model for criteria-aware evaluation. Additionally, we train a writing-enhanced model to demonstrate the effectiveness of our framework’s data curation capability.
3.1 Benchmark Construction
WritingBench is constructed following a delicate pipeline (see Figure 2) that combines model-generated query refinement and human annotation, ensuring diversity and real-world relevance. The two-phrase construction process is illustrated below.
3.1.1 Model-Augmented Query Generation
This phase leverages LLMs to generate an initial set of writing queries, which are enriched and diversified through systematic guidance, with material suggestions provided as needed.
Phase 1: Initial Query Generation
We begin by constructing a two-tiered domain pool grounded in real-world writing scenarios, consisting of 6 primary domains and 100 secondary subdomains (see Figure 3 and Appendix A for detailed descriptions). The selected domains are designed to capture both traditional and emerging user needs for AI-assisted writing, encompassing categories such as Academic & Engineering, Finance & Business, Politics & Law, Literature & Art, Education, Advertising & Marketing. Leveraging the domain and subdomain tags, we prompt two different LLMs (ChatGPT-4o-latest and Claude-3.5-Sonnet) to generate initial writing queries that simulate realistic user requests to maximize diversity (see Appendix B.1).
Phase 2: Query Diversification
To improve the diversity and practical applicability of queries, while taking into account real-world needs, we design a set of query diversification strategies (see Appendix B.2) inspired by Xu et al. (2024), include three core requirements and three auxiliary dimensions:
$\bullet$ Stylistic adjustments (e.g., “Use a friendly and simple tone that kids can easily understand.”)
$\bullet$ Format specifications (e.g., “Follow the IEEE conference template”)
$\bullet$ Length constraints (e.g., “Generate a 500-word executive summary”)
$\bullet$ Personalization (e.g., “From the viewpoint of an educator with two decades of experience.”)
$\bullet$ Content specificity (e.g., “Detail the 2023 Q3 financial metrics”)
$\bullet$ Expression (e.g., Modify the query expression to be shorter.)
The first three categories represent typical writing capabilities, which we will evaluate through specialized assessments. While refining the queries, the LLM simultaneously provides material requirements (e.g., requesting financial reports as input for market analysis queries) (see Appendix B.3). This approach results in enriched queries accompanied by corresponding recommended reference materials.
@@ 3.1.2 Human-in-the-Loop Refinement
This phase incorporates human expertise to verify model-generated queries and supplement material requirements, thereby ensuring their alignment with real-world applications and avoiding harmful content.
Phase 1: Material Collection
At this stage, we engage 30 trained annotators, each compensated at \$18 per hour and possessing specialized expertise proven through domain-knowledge tests. Their role is to collect necessary open-source materials (e.g., public financial statements or legal templates), guided by material requirements generated by LLMs. To minimize errors arising from parsing documents in diverse formats, the annotators carefully extract and verify the most pertinent text segments.
Phase 2: Expert Screening & Optimization
Next, five experts with LLM experience or industry expertise perform data screening. The experts conduct a delicate two-stage filtering process: (1) query adaptation: ambiguous or unrealistic queries are revised to better align with the provided materials and practical scenarios (e.g., adjusting a legal opinion query to reference specific clauses from the supplied statutes), (2) material pruning: redundant or irrelevant content is eliminated from the collected materials, ensuring that the context provided for writing tasks remained focused and relevant while guaranteeing the information is free from harmful content.
Finally, we construct WritingBench, a benchmark comprising 1,239 queries categorized using a two-tiered taxonomy. In comparison to existing writing benchmarks summarized in Table 1, WritingBench exhibits notable advantages in terms of the number of instances, domain diversity, requirement coverage, and variability in input lengths. The detailed statistical distribution of WritingBench is shown in Table 2.
@@ 3.2 Query-Dependent Evaluation Framework
Traditional LLM-as-a-judge evaluations typically rely on fixed evaluation criteria derived from general writing assessment conventions Shao et al. (2024); Bai et al. (2024); Que et al. (2024). However, such static criteria exhibit three critical limitations: (1) Domain exhaustiveness: fixed criteria are inadequate in adapting to specialized domains, failing to address unique characteristics found in areas like technical documentation or creative writing; (2) Requirement specificity: such criteria lack the flexibility to encompass specific requirements related to style, format, or length control; and (3) Material dependency: they are insufficient to verify whether responses appropriately utilize the provided reference materials, such as incorporating data points or maintaining narrative continuity.
To address these challenges, we propose a query-dependent evaluation framework that enables dynamic adaptation to diverse writing scenarios. Our approach comprises two phases:
Phase 1: Dynamic Criteria Generation
As illustrated in Figure 4, given a query $q$ in the WritingBench, the LLM is prompted to automatically generate a set of five evaluation criteria, $C_{q}=\{c_{1},\ldots,c_{5}\}$ (opting for five criteria reflects a common approach observed in many contemporary evaluation contexts Shao et al. (2024); Ma et al. (2024); Marco et al. (2024)), using a carefully designed instruction to ensure structural guidance during criteria specification (see Appendix B.4). Each criterion comprises three components: a concise name summarizing the criterion, an extended description elaborating on the evaluation focus, and detailed scoring rubrics, which provide fine-grained quality levels for the respective evaluation dimensions.
Phase 2: Rubric-based Scoring
For each criterion $c_{i}\in C_{q}$, the evaluator independently assigns a score on a 10-point scale to a response $r$, providing both a score and a justification. The final overall score is computed by averaging scores across all criteria. Detailed prompt is provided in Appendix B.5.
Models Avg Languages Domains Requirements
@@
ChatGPT-4o-latest 8.16 8.3 8.1 8.1 8.1 8.2 8.1 8.4 8.1 8.3 8.7 8.2 8.9 8.2 8.3
o1-Preview 8.15 8.1 8.2 8.0 8.1 8.2 8.2 8.4 8.1 8.2 8.6 8.2 8.8 8.2 8.2
Claude-3.5-Sonnet 7.71 7.7 7.7 7.6 7.5 7.6 7.7 7.9 8.0 7.9 8.5 7.7 8.5 7.9 8.0
Gemini-1.5-Pro 7.78 7.8 7.7 7.7 7.5 7.8 7.9 8.0 7.9 7.9 8.6 7.9 8.8 7.9 8.0
Qwen-Max 8.37 8.4 8.3 8.3 8.3 8.4 8.4 8.5 8.4 8.5 8.7 8.4 9.0 8.4 8.5
@@
Deepseek-R1 8.55 8.7 8.5 8.5 8.5 8.6 8.6 8.7 8.6 8.7 8.9 8.6 9.0 8.6 8.7
Deepseek-V3 7.95 8.0 7.9 7.9 7.8 8.0 7.8 8.2 8.0 8.1 8.6 8.0 8.9 8.0 8.2
Mistral-Large-Instruct 7.64 7.6 7.7 7.7 7.6 7.8 7.3 7.9 7.6 7.7 8.2 7.7 8.7 7.7 7.9
Qwen-2.5-72B-Instruct 7.90 8.0 7.9 8.0 7.8 8.1 7.7 8.2 7.8 8.0 8.3 8.0 8.8 7.9 8.0
Qwen-2.5-7B-Instruct 7.43 7.3 7.5 7.7 7.4 7.6 6.9 7.8 7.3 7.5 7.9 7.6 8.6 7.4 7.5
Llama-3.3-70B-Instruct 7.01 6.7 7.3 7.0 6.9 7.0 6.8 7.3 7.3 7.1 7.8 7.1 8.2 7.0 7.2
Llama-3.1-8B-Instruct 6.35 5.7 6.9 6.6 6.4 6.1 6.0 6.7 6.6 6.4 7.0 6.4 7.6 6.3 6.4

### v4

@@ Abstract
Recent advancements in large language models (LLMs) have significantly enhanced text generation capabilities, yet evaluating their performance in generative writing remains a challenge. Existing benchmarks primarily focus on generic text generation or limited in writing tasks, failing to capture the diverse requirements of high-quality written contents across various domains. To bridge this gap, we present WritingBench, a comprehensive benchmark designed to evaluate LLMs across 6 core writing domains and 100 subdomains.We further propose a query-dependent evaluation framework that empowers LLMs to dynamically generate instance-specific assessment criteria. This framework is complemented by a fine-tuned critic model for criteria-aware scoring, enabling evaluations in style, format and length. The framework’s validity is further demonstrated by its data curation capability, which enables a 7B-parameter model to outperform the performance of GPT-4o in writing. We open-source the benchmark, along with evaluation tools and modular framework components, to advance the development of LLMs in writing.
GitHub Repo: https://github.com/X-PLUG/WritingBench
Leaderboard: https://huggingface.co/spaces/WritingBench/WritingBench
@@ 1 Introduction
Figure 1: Query example with color-coded requirements. Red phrases correlate with gray-shaded writing support materials.
In recent years, LLMs (r1; 4ol) have revolutionized text generation, demonstrating impressive performance across diverse applications, from generating creative content (theatre-human; ex3) and assisting in education (education-survey1; education-survey2) to enhancing professional workflows (storm; metareview). However, generative writing, which requires high levels of coherence, creativity, logical reasoning, and stylistic precision, poses a unique challenge that existing evaluation methods fail to address adequately.
@@
Current evaluation benchmarks for generative writing suffer from two major limitations: 1) Limited scope and diversity in task formulation; and 2) Inadequate evaluation metrics for complex writing tasks. First, there is a significant lack of specialized benchmarks that cover a broad range of writing scenarios. Most existing writing-oriented benchmarks are restricted to single domains, such as fictions (nocha; confederacy-human), and their task formulations tend to be simplistic—often relying on single-sentence prompts (longwriter) or a small set of instruction templates (eqbench; hellobench). Additionally, many benchmarks use homogeneous input materials (hellobench; nocha), limiting their ability to accommodate the complex and customized requirements inherent in real-world writing. As a result, they fail to capture the diversity and intricacies of practical writing tasks (see Figure 1). Second, current automatic evaluation metrics lack the robustness needed for a comprehensive and nuanced assessment of writing quality. While LLM-based evaluation methods show promise in capturing semantic meanings (storm; hellobench; longwriter), they typically rely on a narrow set of predefined criteria (e.g., fluency and coherence). As LLMs continue to evolve with increasingly sophisticated writing capabilities, these static evaluation criteria and frameworks are inadequate for assessing the complex, multi-dimensional nature of writing, including creativity, argumentation strength, and domain-specific adherence.
@@
To address these challenges, we introduce WritingBench, a comprehensive benchmark and robust framework for evaluating general-purpose writing. Our approach begins with a carefully designed secondary domain categorization, grounded in real-world writing needs. We develop a four-stage query construction pipeline (illustrated in Figure 2), where LLMs first generate and diversify writing queries, followed by human-driven material collection and optimization. This process ensures a diverse set of writing tasks with broad domain coverage, varied requirements, and integration of heterogeneous source of materials. To enable a more nuanced evaluation, we propose a query-dependent evaluation framework where LLMs dynamically generates five instance-specific criteria, which are then scored by a fine-tuned critic model. Finally, we integrate the framework to filter writing-specific data and train a small-scale model to verify its ability in identifying high-quality writing samples. Our primary contributions are as follows:
@@
  • We present WritingBench, an open-source writing benchmark comprising 1,000 queries across 6 primary domains and 100 subdomains, featuring style, format and length requirements. WritingBench supports extended-context generation with input ranging from tens to thousands of words, addressing real-world diversity. It facilitates systematic evaluation to identify improvement areas and highlights the potential of chain-of-thought (CoT) processes in creative tasks.
@@
  • We propose a query-dependent evaluation framework that integrates instance-specific criteria generation with a criteria-aware scoring model. It achieves 84% human alignment, significantly surpassing static-criteria baselines (67%, 58%). The effectiveness is further evidenced by its data curation capability-models trained with framework-filtered data outperform GPT-4o in writing.
@@
  • We publicly release WritingBench, including its evaluation protocols, criteria generation tools with an integrated critic model, and writing-enhanced models, to foster further research. Available at: https://github.com/X-PLUG/WritingBench.
@@ 2.1 Writing Benchmarks
Existing writing benchmarks suffer from significant limitations in domain coverage and task granularity. For instance, EQ-Bench (referring to its creative writing subset) encompasses templated queries for story-related tasks (eqbench), while LongBench-Write incorporates length constraints in 120 queries (longwriter); however, they both lack hierarchical domain taxonomies and multi-dimensional requirement specifications (e.g., style and format). Furthermore, most benchmarks rely on fixed instruction templates, short contexts, or materials predominantly from a single source (hellobench; nocha; pros-human-5), rendering them insufficient for addressing the complexity of real-world needs. In contrast, our proposed benchmark fills these gaps by introducing 1,000 free-form queries distributed across 6 primary domains and 100 subdomains, with potential controls over style, format, and length, paired with inputs ranging from tens to thousands of words.
Table 1: Comparison of existing writing benchmarks.
@@
WritingBench 1,000 6 100 1,699 19,361
@@ 2.2 Evaluation Methods
Using LLMs as evaluators has become a prevalent approach for evaluating the quality of generated responses. Typically, researchers pre-define a fixed set of evaluation dimensions applicable across all test instances (confederacy-human; pros-human-5). For example, SuperCLUE (superclue) employs three dimensions, whereas LongBench-Write (longwriter) adopts six dimensions. HelloBench (hellobench) introduces task-specific dimensions, but the dimensions remain consistent across all queries of a given task. Although the LLM-as-a-judge approach enhances scalability, static evaluation dimensions often fail to accommodate the diversity of writing styles and specifications, especially when dealing with inputs with enriched materials. To address this limitation, recent work (fennec) trains a model to dynamically generate evaluation dimensions for individual queries. However, the dimensions remains confined to a small predefined set. In contrast, our query-dependent evaluation framework leverages LLMs to generate diverse and instance-specific criteria while fine-tuning a dedicated critic model to perform the evaluation.
@@ 2.3 Writing-Enhanced Models
Although existing LLMs demonstrate exceptional writing capabilities, researchers strive to further enhance their overall writing proficiency. Recent models, such as Weaver (weaver) and Suri (suri), have exhibited notable domain-specific strengths. For instance, Weaver benefits from over 200B parameter pretraining, supporting four distinct writing domains, while LongWriter (longwriter) specializes in length constraints. However, these models experience substantial performance degradation in cross-domain scenarios and multi-constraint tasks. In this work, leveraging the effectiveness of our evaluation framework, we introduce writing-enhanced models trained on high-quality data, achieving high performance across various tasks.
@@ 3 WritingBench
In this section, we will mainly introduce the four-stage human-AI collaborative construction process of WritingBench, and the query-dependent evaluation framework with its accompanied critic model for criteria-aware evaluation. Additionally, we train a writing-enhanced model to demonstrate the effectiveness of the framework’s data curation capability.
3.1 Benchmark Construction
WritingBench is constructed following a delicate pipeline (see Figure 2) that combines model-generated query refinement and human annotation, ensuring diversity and real-world relevance. The construction process is illustrated below.
3.1.1 Model-Augmented Query Generation
This phase leverages LLMs to generate an initial set of writing queries, which are then enriched and diversified through systematic guidance, with material suggestions provided as needed.
Figure 3: Domain categories in WritingBench. Inner represents the 6 primary domains and outer depicts 100 secondary subdomains (Sub = subdomains per category).
Table 2: Data statistics for WritingBench.
@@ 3.1.1 Model-Augmented Query Generation
Academic & Engineering 167 1,944 15,534
Finance & Business 210 1,867 19,361
Politics & Law 201 2,363 18,317
Literature & Arts 183 1,266 7,675
Education 111 1,345 10,737
Advertising & Marketing 128 984 6,504
@@
Style 442 1,728 19,361
Format 498 1,715 15,534
Length 222 1,362 14,097
@@
0 - 1K 550 470 994
1K - 3K 292 1,832 2,991
3K - 5K 87 3,828 4,966
Over 5K 71 8,053 19,361
Phase 1 - Initial Query Generation: We begin by constructing a two-tiered domain pool grounded in real-world writing scenarios, consisting of 6 primary domains and 100 subdomains (see Figure 3 and Appendix B for detailed domain statistics and descriptions; the construction process of domain system is described in Appendix A.1). These domains are designed to capture both traditional and emerging user needs for AI-assisted writing, categorized by topic and functionality into 6 primary domains: Academic & Engineering, Finance & Business, Politics & Law, Literature & Art, Education, and Advertising & Marketing. Leveraging the primary domain and subdomain tags, we prompt two different LLMs (GPT-4o and Claude-3.5-Sonnet claude-3.5)to produce an extensive pool of initial writing query drafts that simulate realistic user requests to maximize diversity (see Appendix C.2).
Phase 2 - Query Diversification: To enhance the diversity and practical applicability of queries while addressing real-world needs, we design a set of query diversification strategies (see Appendix C.3), inspired by wizardlm. These strategies are divided into three core requirements and three auxiliary dimensions. The core requirements focus on fundamental aspects of writing: (1) Stylistic adjustments (e.g., "Use a friendly and simple tone that kids can easily understand"), (2) Format specifications (e.g., "Follow the IEEE conference template"), and (3) Length constraints (e.g., "Generate a 500-word executive summary"), which will be evaluated through specialized assessments. The auxiliary dimensions address additional considerations: (4) Personalization (e.g., "From the viewpoint of an educator with two decades of experience"), (5) Content specificity (e.g., "Detail the 2023 Q3 financial metrics"), and (6) Expression (e.g., "Modify the query expression to be shorter"). While refining the queries, the LLM simultaneously provides necessary material suggestions (e.g., requesting financial reports as input for market analysis queries) (see Appendix C.4). This structured approach ensures that queries are both diverse and practical.
@@ 3.1.2 Human-in-the-Loop Refinement
The above outputs serve solely to provide a large, linguistically diverse pool of drafts. Human experts then verify these queries and supplement the required materials, ensuring alignment with real-world applications and avoiding harmful content.
Phase 1 - Material Collection: At this stage, we engage 30 trained annotators, compensated at \$18 per hour and possessing specialized expertise proven through domain-knowledge tests. Their role is to collect open-source reference materials needed for the writing tasks (e.g., public financial statements or legal templates). Annotators may either adopt the sources suggested by the LLM-generated material requirements or, when more appropriate, independently identify alternative documents. Each query can be paired with multiple material sources, whereas in some cases no external material is required. To minimize errors caused by parsing documents in varied formats and to eliminate noise from irrelevant content such as web advertisements, the annotators carefully extract and verify only the most pertinent text segments.
Phase 2 - Expert Optimization: Next, five experts with LLM experience or relevant industry expertise conduct data screening. The experts first exclude low-quality data that are unrealistic, overly generic, or potentially harmful. For the remaining entries, a rigorous two-stage review is employed: (1) query adaptation: ambiguous expressions are revised to better align with the provided materials and practical scenarios (e.g., adjusting a legal opinion query to reference specific clauses from the supplied statutes), (2) material pruning: redundant or irrelevant information is eliminated from the accompanying materials, ensuring that the context for writing tasks remains focused, relevant, and devoid of harmful content. Furthermore, experts are encouraged to create original, high-quality queries that are cross-validated by their peers, to further enhance diversity and realism.
Finally, we construct WritingBench, a benchmark comprising 1,000 queries categorized using a two-tiered taxonomy, including 445 queries in Chinese and 555 queries in English. In comparison to existing writing benchmarks summarized in Table 1, WritingBench exhibits notable advantages in terms of the number of instances, domain diversity, requirement coverage, and variability in input lengths. The detailed statistical distribution of WritingBench is shown in Table 3.1.1.
@@ 3.2 Query-Dependent Evaluation Framework
Traditional LLM-as-a-judge evaluations typically rely on fixed evaluation criteria derived from general writing assessment conventions (storm; longwriter; hellobench). However, such static criteria exhibit three critical limitations: (1) Domain exhaustiveness: fixed criteria are inadequate in adapting to specialized domains, failing to address unique characteristics found in areas like technical documentation or creative writing; (2) Requirement specificity: such criteria lack the flexibility to encompass specific requirements related to style, format, or length control; and (3) Material dependency: they are insufficient to verify whether responses appropriately utilize the provided reference materials, such as incorporating data points or maintaining narrative continuity. To address these challenges, we propose a query-dependent evaluation framework that enables dynamic adaptation to diverse writing scenarios. This approach comprises two phases:
Phase 1: Dynamic Criteria Generation: As illustrated in Figure 4, given a query $q$ in the WritingBench, the LLM is prompted to automatically generate a set of five evaluation criteria, $C_{q}=\{c_{1},\ldots,c_{5}\}$ (opting for five criteria reflects a common approach observed in many contemporary evaluation contexts (storm; mops; pros-human-5)), using a carefully designed instruction to ensure structural guidance during criteria specification (see Appendix C.5). We utilize Claude-3.7 for generation, as it demonstrates superior diversity and comprehensiveness in criteria generation compared to models such as GPT-4o(see Appendix A.5.2 for details). Each criterion comprises three components: a concise name summarizing the criterion, an extended description elaborating on the evaluation focus, and detailed scoring rubrics, which provide fine-grained quality levels for the respective evaluation dimensions. The generated criteria are further reviewed by human annotators to confirm their reasonableness and ensure no harmful content.
Table 3: WritingBench performance of LLMs across 6 domains and 3 core requirements evaluated with our critic model (scale: 1-10). The standard deviation is computed over 3 samples. Domains include: (D1) Academic & Engineering, (D2) Finance & Business, (D3) Politics & Law, (D4) Literature & Art, (D5) Education, and (D6) Advertising & Marketing. The writing requirements assessed are: (R1) Style, (R2) Format, and (R3) Length. Here, “C" indicates category-specific scores. The latest results are available on the online leaderboard.
Models Overall Languages Domains Requirements
@@
Claude-3.7-thinking 7.91±0.111 7.9 7.9 7.9 7.8 7.8 8.0 8.0 8.1 7.9 8.7 8.0 8.4 8.0 8.1
Claude-3.7 7.85±0.110 7.9 7.8 7.8 7.8 7.7 7.9 8.0 8.1 7.9 8.6 7.9 8.3 8.0 8.1
Qwen-Max 7.16±0.041 7.2 7.1 7.1 6.9 7.0 7.3 7.4 7.5 7.2 8.3 7.3 7.8 7.2 7.5
o1-Preview 6.89±0.039 6.8 7.0 6.9 6.8 6.7 7.0 7.1 7.2 6.9 8.0 7.0 7.5 7.1 7.3
GPT-4o 6.81±0.028 6.9 6.7 6.8 6.6 6.7 6.8 7.0 7.1 6.9 8.0 7.0 7.5 6.8 6.8
Gemini-1.5-Pro 6.21±0.018 6.2 6.2 6.2 5.8 6.0 6.4 6.6 6.7 6.2 7.2 6.4 7.1 6.4 6.0
@@
Deepseek-R1 7.70±0.053 8.0 7.5 7.6 7.4 7.6 7.8 7.8 8.1 7.7 8.4 7.9 8.3 7.7 7.5
Deepseek-V3 6.35±0.022 6.3 6.4 6.4 6.1 6.2 6.3 6.6 6.8 6.4 7.6 6.5 7.1 6.5 6.4
Mistral-Large-Instruct 6.00±0.076 5.9 6.1 6.2 5.9 5.9 5.7 6.4 6.4 6.1 7.3 6.1 6.5 6.0 6.0
Qwen-2.5-72B-Instruct 6.40±0.061 6.4 6.4 6.6 6.2 6.4 6.2 6.7 6.6 6.5 7.7 6.5 6.9 6.5 6.5
Qwen-2.5-7B-Instruct 5.64±0.083 5.5 5.8 5.9 5.6 5.6 5.1 6.1 5.9 5.7 7.0 5.7 6.1 5.6 5.6
Llama-3.3-70B-Instruct 5.05±0.011 4.5 5.5 5.1 4.9 4.8 4.8 5.3 5.6 5.0 5.0 5.1 5.9 5.1 5.0
Llama-3.1-8B-Instruct 4.42±0.004 3.7 5.0 4.1 4.4 4.0 4.1 4.7 5.0 4.4 4.4 4.5 5.3 4.4 4.3